# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

### v1

@@ 3. Reanalysis: Methods
Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model consists of model inputs $C$ and $H$ (the number of COVID cases and hospitalizations in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination). These model inputs are combined to project the benefit of vaccinating the test population, and this projected benefit is compared to the model’s corresponding risk inputs, the number of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose. This comparison, or output of the model, depends on what values are used for each of the model inputs. The FDA performed the comparison under six different scenarios, each corresponding to a particular choice of input values.
Because the FDA relied on CDC data to supply projections of values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH.
We extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination, focusing on the hospitalizations endpoint: those prevented by vaccination vs. those caused by VAM/P. Our model has the same structure as the FDA’s but has infections (not reported cases) as a key model input and links infections to hospitalizations via another key input, the infection hospitalization rate IHR. Additional inputs account for the fraction of the test population having a prior COVID infection as well as for the protection against infection and hospitalization provided by vaccination, by prior-infection, or by both (hybrid protection). Model inputs are combined to project the hospitalizations-prevented benefit of vaccinating a test population, and this projected benefit is compared to the model’s risk input, the projected number of VAM/P hospitalizations expected to occur in the course of achieving a fully vaccinated test population. In our modeling, ”infection” is any COVID infection, symptomatic or asymptomatic.
Test population. Our model is based on a hypothetical population of 1 million males 18–25, so that our test population consists of those at highest risk of VAM/P (among the groups considered in the FDA’s mRNA-1273 assessment). We assume that the test population is representative, as of 1 January 2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection. We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22). The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population. We call a test population consisting of infection naïve unvaccinated persons a baseline test population.
We begin by introducing the inputs to our model.
@@ 3.1. Model Inputs
Infection rate $I_{r}$. Rather than base our projections of COVID prevalence on reported cases, we base our modeling on $I_{r}$, the fraction of a baseline test population of unvaccinated infection-naïve males 18–25 projected to become infected (symptomatic and asymptomatic) over the FDA’s 5-month evaluation period.
Infection-hospitalization rate $H_{r}$. We translate infections to projected hospitalizations using an estimated infection-hospitalization rate $H_{r}$ for unvaccinated, infection-naïve males 18–25 years old. $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization.
Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection before the start of the 5-month evaluation period.
Vaccine effectiveness against infection $E_{v}$. This is the fraction by which the baseline infection rate $I_{r}$ is reduced due to the protection afforded by vaccination. That is, the infection rate for a fully vaccinated test population is $I_{r}(1-E_{v})$.
@@
VAM/P hospitalizations $H_{\rm VAM/P}$. The final input to our model is $H_{\rm VAM/P}$, which is the projected number of VAM/P hospitalizations expected to occur in the course of attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males. To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
@@ 3.2. Combining Model Inputs to Produce Outputs
Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1/1/22 – 5/31/22. Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period. Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period. The difference $U-V$ is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section.
In our modeling, we consider only the possibility of hospitalizations associated with test-population members’ first infections during the evaluation period. We expect second infections over the course of a five-month period to be rare; moreover, infection-hospitalization risk will drop significantly after infection [16, 17, 18].121212 The meta-analyses [22, 23] suggest prior-infection provides greater protection against Omicron hospitalization than full vaccination among the COVID-naïve—see Section 7 of Supplement 1 for details. Our decision to analyze hospitalization risk associated with only first infections during the 5-month evaluation period allows us to assume, as did the FDA, that all effectiveness factors contributing to the analysis remain constant over the evaluation period.
Deriving $U$. To derive $U$, we will assume that the test population remains unvaccinated during the evaluation period. As a first step, suppose that the entire test population is infection naïve. Then, in this baseline population, we expect $(1\ {\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1\ {\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $(1\ {\rm m})$ is an abbreviation for (1 million).
@@
Deriving $V$. Next we assume that every individual in the test population is fully vaccinated. Deriving the total number of Omicron hospitalizations $V$ for this population is analogous to deriving $U$, but more factors come into play because we must account for the protective effects of both vaccination and prior infection. Since all members of this test population are vaccinated, those with prior infection have hybrid protection. Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals. The first stage depends on the model inputs $E_{v}$ and $E_{h}$, while the second stage uses $HRR_{v}$ and $HRR_{h}$. Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
Then in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection. So the total number of hospitalizations in the population is
@@
Hospitalizations prevented by vaccination. The difference between total Omicron hospitalizations in an unvaccinated test population and total hospitalizations in a fully vaccinated test population is the hospitalizations-prevented benefit of vaccination for the test population:
@@
We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent to assuming in our model that all persons in the test population are infection-naïve, i.e. that $F_{pi}=0$. Notice that if $F_{pi}=0$ in Equation (3.2.3), then its right-hand side reduces to
@@
[…]
Applying the model for benefits-risk assessment of vaccination. For the model to output a meaningful assessment of the benefits and risks of vaccination, we must make good estimates for the model inputs appearing in (3.2.3), as well as for the remaining model input: $H_{\rm VAM/P}$, which by itself describes the risk side of the assessment. Then the hospitalizations-prevented benefit $U-V$ can be compared to $H_{\rm VAM/P}$ to complete the analysis.
@@ 3.3. Estimating Values for Model Inputs
Like the FDA, we adopt a scenario-based structure for our benefit-risk assessment to account for uncertainties associated with pandemic dynamics as well as with the effectiveness of vaccine and prior-infection protection. Table 1 of Section 2 above summarizes the model-input assumptions the FDA used for each of the six scenarios of its mRNA-1273 assessment [9]. Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario. Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’ s Scenarios 1–3 in that our COVID incidence assumptions are derived from the FDA’s corresponding assumptions. Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other input variables represent our best estimates. Thus, we consider Scenario D to be our most likely scenario. In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness. Model inputs and outputs for our scenarios are described in Table 6 of the next section, while the methods we use to estimate values of model inputs are summarized below. Our estimates are based on data available to the FDA before 1/$2$
Estimating $I_{r}$. As we noted in Section 2.2, the FDA’s COVID-case rate assumption for its most likely scenario, Scenario 1, is so large that it corresponds to the assumption that essentially all unprotected (i.e., unvaccinated) 18–25-year-old males become infected over the 5-month evaluation period. Thus, in Scenario A, our reanalysis of Scenario 1, we too assume all unprotected persons in the test population become infected. However, in our modeling, the unprotected are the infection-naïve unvaccinated. Assuming all such persons become infected over 5 months is equivalent to setting $I_{r}=1$, which is our Scenario-A value for $I_{r}$. As discussed in Section 1 of Supplement 1, for our Scenarios B and C, values of $I_{r}$ are determined by the FDA’s COVID-incidence assumptions for its Scenarios 2 and 3, respectively, while for our Scenarios D and E, $I_{r}$ is derived from the assumption that over the evaluation period the increase in COVID-infection level among 18–25-year-old males is twice that of the general population during the second COVID wave in the U.S. (10/1/20–2/28/21).
Estimating $H_{r}$. Our estimate for $H_{r}$ is derived from an IHR model [51] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [52]. Applying the model, we obtain a “with COVID” IHR of 0.37% for the infection-naïve unvaccinated of ages 18–25 and reduce this rate assuming 40% of hospitalizations attributed to COVID are incidental, arriving at an IHR of $0.22\%\approx 0.6\cdot 0.37\%$ (for infection-naïve unvaccinated 18–25 year-olds). The 0.22% rate for the general population of 18–25 year-olds in the U.S. (males and females) is consistent with U.S. military COVID-19 hospitalization data from the period 1 February 2020 through 4 January 2021 [53]. See Supplement 1 for further details concerning our derivation of the IHR value $0.22\%$.
In its benefit-risk assessments [8] and [9] for, respectively, Pfizer’s and Moderna’s mRNA vaccines, the FDA relied on sex-specific CDC hospitalizations data. The case and hospitalization rates per 100,000 persons in Table 2 of [8] suggests that for the 18–24 age range, females are 2.53 times as likely as males to be hospitalized with COVID-19 and 1.21 times as likely to have a case of COVID-19.131313Moreover, these relationships are consistent with those of [54, Table 1] based on data from 3/1/20 to 3/1/22. In Section 2 of Supplement 1, we use hospitalization and case data from Table 2 of [8] to estimate the IHR for unvaccinated, infection-naïve males 18–25 to be $0.14\%$ (and for females 18–25 to be $0.29\%$).
Estimating $F_{pi}$. For our Scenarios A–C we use the value $F_{pi}=0.62$, which is the average of (a) the CDC’s estimate, mentioned earlier, that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [26] with (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1 January 2022. For Scenarios D and E, we use $F_{pi}=0.69$. These values of $F_{pi}$ may be underestimates because they rely on the CDC’s infected rate of 54.9% by 10/1/21 for the general population of 18–49 year-olds, and we would expect a somewhat higher infected rate among the unvaccinated.
Estimating $E_{v}$, $HRR_{v}$. Recall that the FDA used VE = 0.30 and VEH = 0.72 for all its Omicron-based scenarios [9, Table 1]. For reasons discussed in Supplement 1, for all our scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$. Hence, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the VEH used by FDA. We remark that the FDA’s choices of $VEH=0.72$ and $E_{v}=0.30$ are not consistent with the data and methods of the UK briefing [34] from which the FDA obtained its VEH estimate.
Estimating $E_{pi}$ and $E_{h}$. Our values $E_{pi}=0.47$ and $E_{h}=0.57$ are derived (see Supplement 1) from a December 2021 study [27], based on data from South Africa. [^14] As discussed in Supplement 1, our effectiveness estimates involving prior infection are consistent with those of meta-analyses [22] and [23].
[…]
{adjustwidth}
-1cm
Table 5. Summary of Model Inputs
@@
$I_{r}$: Baseline-infection rate—fraction of an unvaccinated, infection-naïve test-population of 18–25-year-old males expected to contract an infection over the 5-month evaluation period. Infections need not accumulate at a constant rate $0\leq I_{r}\leq 1$ $1$
$H_{r}$: Baseline infection-hospitalization rate for infection-naïve unvaccinated males 18–25 $0\leq H_{r}\leq 0.01$ $0.0014$
@@
*Remarks. (i) All measures of effectiveness of protection against infection ($E_{pi}$, $E_{v}$, and $E_{h}$) are relative to infection-naïve, unvaccinated persons. (ii) All hospitalization-risk reductions ($HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$) are relative to unvaccinated persons experiencing primary infections.

### v3

@@ 3. Reanalysis: Methods
Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model has inputs $C$ and $H$ (the numbers of evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination). These model inputs are combined to project the cases-prevented and hospitalizations-prevented benefits of vaccinating the test population, and these benefits are compared to the model’s corresponding risk inputs, the numbers of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
Our model, which has hospitalizations as its endpoint, is algebraic (like the FDA’s) and thus is widely accessible. Our model extends the FDA’s by not only incorporating prior-infection protection but also by being essentially equivalent to the FDA’s model if prior infection is assumed to provide no protection. Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH. In contrast, our model does provide these links. For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection hospitalization rate IHR. Additional input variables are described in Section 3.1 below and summarized in Table 5.
Model inputs are combined to project the hospitalizations-prevented benefit of vaccinating a test population, and this projected benefit is compared to the model’s risk input, the projected number of VAM/P hospitalizations expected to occur in the course of achieving a fully vaccinated test population. In our modeling, ”infection” is any COVID infection, symptomatic or asymptomatic.
Test population. Our model is based on a hypothetical test population of 1 million males 18–25. We assume that the test population is representative, as of 1/1/2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection. We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22). The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population. We call a test population consisting of infection naïve unvaccinated persons a baseline test population.
@@ 3.1. Model Inputs
Infection rate $I_{r}$. We base our modeling on $I_{r}$, the fraction of a baseline test population of unvaccinated infection-naïve males 18–25 projected to become infected (symptomatic and asymptomatic) over the FDA’s 5-month evaluation period.
Infection-hospitalization rate $H_{r}$. We translate infections to projected hospitalizations using an estimated infection-hospitalization rate $H_{r}$ for unvaccinated, infection-naïve males 18–25 years old. $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization. We emphasize that $H_{r}$ is a sex-specific rate. The FDA provides data in Table 2 of [10] indicating that for the 18–24 age-range $H_{r}$ for males is significantly less that that for females. [^7] The data of Table 1 of [54] is consistent with that of [10, Table 2].
Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection prior to the evaluation period.
Vaccine effectiveness against infection $E_{v}$. This is the fraction by which the baseline infection rate $I_{r}$ is reduced due to the protection afforded by vaccination; i.e., the infection rate for a fully vaccinated test population is $I_{r}(1-E_{v})$.
@@
VAM/P hospitalizations $H_{\rm VAM/P}$. Our final model input $H_{\rm VAM/P}$ is the projected number of VAM/P hospitalizations expected to occur in attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males. To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations. In Section 2.3, based on data available before 1/1/22, we obtained an estimated VAM/P hospitalization rate of 250 per million 2nd doses of mRNA-1273 given to 18–25-year-old males. In Section 6.7 of Supplement 1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose rate of the FDA’s BLA-Memo Scenario—compare the first two rows of Table 4 above. Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$. [^8] In Supplement 1 (Appendix 2, Section 2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 1/22/22, showing the data suggests a VAM/P hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
@@ 3.2. Combining Model Inputs to Produce Outputs
Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1/1/22 – 5/31/22. Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period. Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period. The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, our model is deterministic).
In our modeling, we consider only the possibility of hospitalizations associated with test-population members’ first infections during the evaluation period. We expect second infections over a 5-month period to be rare; moreover, infection-hospitalization risk will drop significantly after infection [15, 16, 17]. Our decision to analyze hospitalization risk associated with only first infections during the 5-month evaluation period allows us to assume, as did the FDA, that all effectiveness factors contributing to the analysis remain constant over the evaluation period. Each effectiveness value represents average effectiveness over the evaluation period; in this way, our modeling accounts for waning of vaccine and prior-infection protection.
Deriving $U$. To derive $U$, we will assume that the test population remains unvaccinated during the evaluation period. As a first step, suppose that the entire test population is infection naïve. Then, in this baseline population, we expect $(1\ {\rm m})\cdot I_{r}$ infections to develop over 5 months, resulting in $(1\ {\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations, where $1\ {\rm m}=$ 1 million.
@@
Deriving $V$. Next we assume that every individual in the test population is fully vaccinated. Since all members of this test population are vaccinated, those with prior infection have hybrid protection. Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals. Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
Then, in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection. So the total number of hospitalizations in the population is
@@
Hospitalizations prevented by vaccination. As noted above, the difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population:
@@
We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e. that $F_{pi}=0$. With $F_{pi}=0$ in Equation (3.2.3), its right-hand side is
(3.2.4) $\left[\text{Expected Baseline Hospitalizations}\right]\left(1-(1-E_{v})(1-HRR_{v})\right),$
@@ 3.3. Estimating Values for Model Inputs
Like the FDA, we account for uncertainties in model inputs by adopting a scenario-based assessment. The FDA’s model-input assumptions for its six scenarios of [11] are summarized in Table 1 of Section 2. Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario. Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s corresponding assumptions. Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other inputs represent our best estimates. Thus, Scenario D is our most likely scenario. In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness.
Table 5 below summarizes model inputs with footnotes providing some comments on the estimation of their values—see Supplement 1 for details.
Table 5. Summary of Input Variables*
@@
$I_{r}$: Baseline-infection rate—fraction of an unvaccinated, infection-naïve test-population of 18–25-year-old males expected to contract an infection over the 5-month evaluation period. Infections need not accumulate at a constant rate $0\leq I_{r}\leq 1$ 11
$H_{r}$: Baseline infection-hospitalization rate for infection-naïve unvaccinated males 18–25 $\leq H_{r}\leq 0.01 0.0$0140.0014
@@
*Remarks. (i) Estimates are based on data available before 1/22/22. (ii) $E_{pi}$, $E_{v}$, and $E_{h}$ are relative to infection-naïve, unvaccinated persons. (iii) $HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$ are relative to unvaccinated persons experiencing primary infections. (iv) Our Scenario-A value $I_{r}=1$ reflects the assumption that all infection-naïve unvaccinated males 18–25 become infected over the evaluation period, consistent with FDA Scenario-1 assumptions (see Section 2.2). (v) We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [56]. We assume a 40% incidental-hospitalization rate and make sex-based adjustments based on [10, Table 2]. (See [Supplement 1, Section 2]) (vi) $F_{pi}=0.62$ is the average of (a) the CDC’s estimate that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [25] and (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1/1/2022. (vii) For reasons discussed in Supplement 1, for all scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$; thus, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the FDA’s VEH (viii) Estimates for $E_{pi},E_{h},HRR_{pi}$, and $HRR_{h}$, are based on data from [26] and [27]. $HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combine with $E_{pi}=0.47$ and $E_{h}=0.57$ to yield overall estimates for protection against hospitalization—88.9% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [21] and [22].